# Symplectic Quantization: numerical results for the Feynman propagator on a 1+1 lattice and the theoretical relation with Quantum Field Theory

### v2

@@ VI Feynman propagator: numerical results
The strategy is very simple: having set the coefficient of the non-linear interaction $\lambda$ to a small but finite value, $\lambda=0.001$, we have run the symplectic dynamics with fringe boundary conditions until stationarity is reached at a certain time, which we call $\tau_{\text{eq}}$. According to the premises of Sec. II, where we assumed that at long enough times the symplectic quantization dynamics allows us to sample an equilibrium ensemble, we have checked that equipartition between positional and kinetic degrees of freedom is in fact reached. In Fig. 4 is shown how, for two given choices of $k=\{k_{0},k_{1}\}$ (corresponding respectively to small and large scales), we have that $\overline{E}_{\text{harm}}(k,\tau)$ and $\overline{E}_{\text{kin}}(k,\tau)$ reach asymptotically a value close to $1/2$, starting respectively from $E_{\text{harm}}(k,0)=0$ and $E_{\text{kin}}(k,0)=1$, where the two energies are defined respectively as
[…]
@@
In Fig. 5 we show (bottom panel) the result for the Fourier spectrum of the two-point correlation function obtained by setting all the parameters of the simulation and the initial conditions as declared at the end of Sec. III, apart from the value of the mass that is set here at $m=3.0$ in order to better appreciate the shape of the propagator, and taking the value $\lambda=0.001$ for the non-linearity parameter. In order to compare our numerical data at small non-linearity with the theory, we have also reported in the top panel of Fig.5 the theoretical shape of the free Feynman propagator $G_{\text{th}}(k_{0},k_{1})$ on a discretized space-time grid in $1+1$ dimensions, using for the lattice the same parameters of the simulation, i.e., $a=1.0$, $m=1.0$, and $M_{\mu}=128$, where $G_{\text{th}}(k_{0},k_{1})$ reads as
@@
Figure 6: Real space two-point correlation function for a $\lambda\phi^{4}$ theory in $1+1$ space-time dimensions with fringe boundary conditions, lattice spacing $a=1.0$, lattice side $M_{\mu}=128$, mass $m=1.0$ and nonlinearity $\lambda=0.001$. Top): exponential decay along the direction parallel to the $x_{1}$ axis; Bottom): oscillations along the direction parallel to the $x_{0}=ct$ axis.
@@
$\Delta_{F}(x-y)=\frac{1}{(2\pi)^{2}}\int d^{2}k\leavevmode\nobreak\ \frac{e^{ik(x-y)}}{k^{2}-m^{2}},$ (59)
@@
$\Delta_{F}(x-y)$ $\sim e^{im|x-y|}\quad\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ x-y\leavevmode\nobreak\ ||\leavevmode\nobreak\ x_{0}$
$\Delta_{F}(x-y)$ $\sim e^{-m|x-y|}\quad\text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ x-y\leavevmode\nobreak\ ||\leavevmode\nobreak\ x_{1}.$ (60)
Clearly, when the same correlation function is sampled on a finite and discrete grid there will be finite-size effects at play so that, for instance, also the oscillations along the time-like direction will be slightly modulated by a tiny exponential decay: this is precisely what we find in numerical simulations. In Fig. [6] are shown, respectively in top and bottom panels, the exponential decay along the purely space-like direction and the oscillations along the purely time-like direction, obtained for the following choice of parameters: $a=1.0$, $M_{\mu}=128$, mass $m=1.0$ and nonlinearity $\lambda=0.001$. Let us notice that the value of the mass which can be obtained from either the fit of the exponential decay as $C(\Delta x_{1})\sim e^{-m\Delta x_{1}}$ or the oscillating part as $C(\Delta x_{0})\sim e^{im\Delta x_{0}}$ is $m\sim 2.06\pm 0.04$, i.e., quite different from the value $m=1$ put in the Lagrangian. This effect, which we do not find for the deterministic dynamics in Euclidean space-time, is most probably a finite-size effect related to nature of the new fringe boundary conditions adopted. We will devote our next effort in the numerical investigation of symplectic quantization to carefully study the finite-size effects induced by fringe boundaries.
VII Canonical form of Minkowskian statistical mechanics
In the previous section we have shown that the Hamiltonian dynamics of a quantum field theory with an additional time paramter $\tau$ and corresponding conjugated momenta allows to recover qualitatively well the shape of the free Feynman propagator for a small value of the interaction constant $\lambda$. It is therefore legitimate to wonder which is the precise relation between the correlation functions obtained in this generalized microncanonical ensemble and the one generated by the Feynman path integral and/or the corresponding Euclidean Field Theory. As a first step in this direction we propose an explicit calculation of the microcanonical partition function in the large-$M$ limit, where $M$ is the number of degrees of freedom. The calculation shows that also for this peculiar system, where the microcanonical ensemble is built on the conservation of an action rather than an energy functional, the sampling of fluctuations in this ensemble is formally equivalent to the sampling of a canonical one at a temperature $\hbar$, namely fields fluctuations are sampled with probability $\exp(S/\hbar)$.
@@ VII Canonical form of Minkowskian statistical mechanics
$\Omega[\mathcal{A},J]=\int\mathcal{D}\phi\mathcal{D}\pi\leavevmode\nobreak\ \delta\left(\mathcal{A}-\mathbb{H}[\phi,\pi]+\int d^{d}x\,J(x)\phi(x)\right).$ (61)
Since we are using the lattice a regularizer for the theory, the field can be conveniently expanded in an orthonormal basis as follows [19]:
@@
[…]
Contrary to the usual convention, there is no $\hbar$ in this measure. In a $d$-dimensional box of volume $L^{d}$ with lattice spacing $a$, the number of basis functions is [19]:
@@
[…]
$=\frac{(2\pi)^{\frac{M}{2}}}{\Gamma\left(\frac{M}{2}\right)}\leavevmode\nobreak\ R^{M-2},$ (69)
@@
$\Omega[\mathcal{A}_{z},J]$ $=\leavevmode\nobreak\ \kappa_{M}\left(\frac{\hbar M}{2z}\right)^{\frac{M}{2}}\sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{z}{\hbar}\right)^{n}\left(\frac{2}{M}\right)^{n}\frac{\Gamma(\frac{M}{2}+1)}{\Gamma(\frac{M}{2}+1-n)}$
×$\times\int d^{d}x_{1}\ldots d^{d}x_{n}\leavevmode\nobreak\ J(x_{1})\ldots J(x_{n})\int\mathcal{D}_{M}\phi\leavevmode\nobreak\ \phi(x_{1})\ldots\phi(x_{n})\left(1+\frac{2}{M}\frac{z}{\hbar}S[\phi]\right)^{\frac{M}{2}-n}\,.$ (73)
@@
$\left(1+\frac{2}{M}\frac{z}{\hbar}S[\phi]\right)^{\frac{M}{2}-n}=e^{\frac{z}{\hbar}S[\phi]}\exp\Bigg{(}\sum_{j=1}^{\infty}(-1)^{j}\leavevmode\nobreak\ \frac{2(2j-2)!!}{(j+1)!}\left(\frac{z}{\hbar}\frac{S[\phi]}{M}\right)^{j}\leavevmode\nobreak\ [jS[\phi]+(j+1)n]\Bigg{)}\,.$ (74)
@@
$\Omega$ $[\mathcal{A}_{z},J]=\kappa_{M}\leavevmode\nobreak\ \left(\frac{\hbar M}{2z}\right)^{\frac{M}{2}}\leavevmode\nobreak\ \sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{z}{\hbar}\right)^{n}\left(\frac{2}{M}\right)^{n}\frac{\Gamma(\frac{M}{2}+1)}{\Gamma(\frac{M}{2}+1-n)}$
×$\times\int d^{d}x_{1}\ldots d^{d}x_{n}\leavevmode\nobreak\ J(x_{1})\ldots J(x_{n})\leavevmode\nobreak\ \left\langle\phi(x_{1})\ldots\phi(x_{n})\leavevmode\nobreak\ e^{\sum_{j=1}^{\infty}(-1)^{j}\leavevmode\nobreak\ \frac{2(2j-2)!!}{(j+1)!}\left(\frac{z}{\hbar}\frac{S[\phi]}{M}\right)^{j}\leavevmode\nobreak\ [jS[\phi]+(j+1)n]}\right\rangle,$
[…]
$\left\langle\phi(x_{1})\ldots\phi(x_{n})\leavevmode\nobreak\ e^{\sum_{j=1}^{\infty}(-1)^{j}\leavevmode\nobreak\ \frac{2(2j-2)!!}{(j+1)!}\left(\frac{z}{\hbar}\frac{S[\phi]}{M}\right)^{j}\leavevmode\nobreak\ [jS[\phi]+(j+1)n]}\right\rangle=\left\langle\phi(x_{1})\ldots\phi(x_{n})\right\rangle+\sum_{j=1}^{\infty}\frac{c_{j}(M,n)}{M^{j}}\left\langle\phi_{1}(x_{1})\ldots\phi_{n}(x_{n})\leavevmode\nobreak\ S^{j}[\phi]\right\rangle,$ (75)
@@
$\lim_{M\to\infty}\left\langle\phi_{1}(x_{1})\ldots,\phi_{n}(x_{n})\frac{c_{j}(M,n)}{M^{j}}S^{j}[\phi]\right\rangle=0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ j,$ (78)
@@
$\Omega[\mathcal{A}_{z},J]$ $=\kappa_{M}\leavevmode\nobreak\ \left(\frac{\hbar M}{2z}\right)^{\frac{M}{2}}\leavevmode\nobreak\ \sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{z}{\hbar}\right)^{n}\leavevmode\nobreak\ \int d^{d}x_{1}\ldots d^{d}x_{n}\leavevmode\nobreak\ J(x_{1})\ldots J(x_{n})\leavevmode\nobreak\ \left\langle\phi(x_{1})\ldots\phi(x_{n})\right\rangle$
$=\kappa_{M}\leavevmode\nobreak\ \left(\frac{\hbar M}{2z}\right)^{\frac{M}{2}}\int\mathcal{D}\phi_{M}\leavevmode\nobreak\ e^{\frac{z}{\hbar}S[\phi]}\leavevmode\nobreak\ \sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{z}{\hbar}\right)^{n}\leavevmode\nobreak\ \int d^{d}x_{1}\ldots d^{d}x_{n}\leavevmode\nobreak\ J(x_{1})\ldots J(x_{n})\leavevmode\nobreak\ \phi(x_{1})\ldots\phi(x_{n})$
$=\kappa_{M}\leavevmode\nobreak\ \left(\frac{\hbar M}{2z}\right)^{\frac{M}{2}}\int\mathcal{D}\phi_{M}\leavevmode\nobreak\ e^{\frac{z}{\hbar}S[\phi]+\frac{z}{\hbar}\int d^{d}x\leavevmode\nobreak\ J(x)\phi(x)}\,.$ (79)
@@
This is completely in agreement with the results of numerical simulations in the microcanonical ensemble, where the Hamiltonian dynamics of the free theory develops run-away solutions. Second, the result of our derivation in Eq. (80) shows us that this new ”canonical Minkowskian measure” can be connected to the standard Feynman path integral by means of analytic continuation in the dimensionless parameter $z$. More investigations in this direction are actually in progress.
VIII Conclusions and Perspectives

### v3

@@ 6 Feynman propagator: numerical results
The strategy is very simple: having set the coefficient of the non-linear interaction $\lambda$ to a small but finite value, $\lambda=0.001$, we have run the symplectic dynamics with fringe boundary conditions until stationarity is reached at a certain time, which we call $\tau_{\text{eq}}$. According to the premises of Sec. 2, where we assumed that at long enough times the symplectic quantization dynamics allows us to sample an equilibrium ensemble, we have checked that equipartition between positional and kinetic degrees of freedom is in fact reached. In Fig. 3 is shown how, for two given choices of $k=\{k_{0},k_{1}\}$ (corresponding respectively to small and large scales), we have that $\overline{E}_{\text{harm}}(k,\tau)$ and $\overline{E}_{\text{kin}}(k,\tau)$ reach asymptotically a value close to $1/2$, starting respectively from $E_{\text{harm}}(k,0)=0$ and $E_{\text{kin}}(k,0)=1$, where the two energies are defined respectively as
$\overline{E}_{\text{harm}}(k,\tau)$ $=\frac{1}{\tau}\int_{0}^{\tau}ds~\frac{1}{2}\omega_{k}^{2}|\phi(k,s)|^{2}$
$\overline{E}_{\text{kin}}(k,\tau)$ $=\frac{1}{\tau}\int_{0}^{\tau}ds~\frac{1}{2}|\pi(k,s)|^{2}.$ (56)
@@
In Fig. 4 we show (bottom panel) the result for the Fourier spectrum of the two-point correlation function obtained by setting all the parameters of the simulation and the initial conditions as declared at the end of Sec. 3, apart from the value of the mass that is set here at $m=3.0$ in order to better appreciate the shape of the propagator, and taking the value $\lambda=0.001$ for the non-linearity parameter. In order to compare our numerical data at small non-linearity with the theory, we have also reported in the top panel of Fig.4 the theoretical shape of the free Feynman propagator $G_{\text{th}}(k_{0},k_{1})$ on a discretized space-time grid in $1+1$ dimensions, using for the lattice the same parameters of the simulation, i.e., $a=1.0$, $m=1.0$, and $M_{\mu}=128$, where $G_{\text{th}}(k_{0},k_{1})$ reads as
@@
Figure 5: Real space two-point correlation function for a $\lambda\phi^{4}$ theory in $1+1$ space-time dimensions with fringe boundary conditions, lattice spacing $a=1.0$, lattice side $M_{\mu}=128$, mass $m=1.0$ and nonlinearity $\lambda=0.001$. Top: exponential decay along the direction parallel to the $x_{1}$ axis; bottom: oscillations along the direction parallel to the $x_{0}=ct$ axis.
@@
$\Delta_{F}(x-y)=\frac{1}{(2\pi)^{2}}\int d^{2}k~\frac{e^{ik(x-y)}}{k^{2}-m^{2}},$ (59)
@@
$\Delta_{F}(x-y)$ $\sim e^{im|x-y|}\quad~\text{for}~~~x-y~||~x_{0}$
$\Delta_{F}(x-y)$ $\sim e^{-m|x-y|}\quad\text{for}~~~x-y~||~x_{1}.$ (60)
Clearly, when the same correlation function is sampled on a finite and discrete grid there will be finite-size effects at play so that, for instance, also the oscillations along the time-like direction will be slightly modulated by a tiny exponential decay: this is precisely what we find in numerical simulations. In Fig. [5] are shown, respectively in top and bottom panels, the exponential decay along the purely space-like direction and the oscillations along the purely time-like direction, obtained for the following choice of parameters: $a=1.0$, $M_{\mu}=128$, mass $m=1.0$ and nonlinearity $\lambda=0.001$. Let us notice that the value of the mass which can be obtained from either the fit of the exponential decay as $C(\Delta x_{1})\sim e^{-m\Delta x_{1}}$ or the oscillating part as $C(\Delta x_{0})\sim e^{im\Delta x_{0}}$ is $m\sim 2.06\pm 0.04$, i.e., quite different from the value $m=1$ put in the Lagrangian. This effect, which we do not find for the deterministic dynamics in Euclidean space-time, is most probably a finite-size effect related to the propagation of signals across fringe boundary conditions. We made some attempts, discussed in Appendix B, to investigate a possible interplay between the measured mass discrepancy with the way fringe boundary conditions are imposed, but we didn’t find any clear indication on the possible origin of the effect. A stronger effort is for sure necessary to put under control the finite-size effects related to fringe boundary conditions: we plan to devote another paper to this problem.
7 Canonical form of Minkowskian statistical mechanics
In Sec. 6 we have shown that the Hamiltonian dynamics of a quantum field theory with an additional time paramter $\tau$ and corresponding conjugated momenta allows to recover qualitatively well the shape of the free Feynman propagator for a small value of the interaction constant $\lambda$. It is therefore legitimate to wonder which is the precise relation between the correlation functions obtained in this generalized microncanonical ensemble and the one generated by the Feynman path integral and/or the corresponding Euclidean Field Theory. As a first step in this direction we propose an explicit calculation of the microcanonical partition function in the large-$M$ limit, where $M$ is the number of degrees of freedom. The calculation shows that also for this peculiar system, where the microcanonical ensemble is built on the conservation of an action rather than an energy functional, the sampling of fluctuations in this ensemble is formally equivalent to the sampling of a canonical one at a temperature $\hbar$, namely fields fluctuations are sampled with probability $\exp(S/\hbar)$.
@@ 7 Canonical form of Minkowskian statistical mechanics
$\Omega[\mathcal{A},J]=\int\mathcal{D}\phi\mathcal{D}\pi~\delta\left(\mathcal{A}-\mathbb{H}[\phi,\pi]+\int d^{d}x\,J(x)\phi(x)\right).$ (61)
Since we are using the lattice a regularizer for the theory, the field can be conveniently expanded in an orthonormal basis as follows STROMINGER :
@@
$\int\mathcal{D}_{M}\phi\equiv\prod_{n=1}^{M}\int_{-\infty}^{\infty}dc_{n}.$ (64)
Contrary to the usual convention, there is no $\hbar$ in this measure. In a $d$-dimensional box of volume $L^{d}$ with lattice spacing $a$, the number of basis functions is STROMINGER :
@@
$I_{M}(R)$ $=\int_{-\infty}^{\infty}dx_{1}\ldots dx_{M}~\delta\left(\frac{1}{2}\sum_{i=1}^{M}x_{i}^{2}-R^{2}\right)=\frac{(2\pi)^{\frac{M}{2}}}{\Gamma\left(\frac{M}{2}\right)}~R^{M-2},$ (68)
@@
$\Omega[\mathcal{A}_{z},J]=$ $~\kappa_{M}\left(\frac{\hbar M}{2z}\right)^{\frac{M}{2}}\sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{z}{\hbar}\right)^{n}\left(\frac{2}{M}\right)^{n}\frac{\Gamma(\frac{M}{2}+1)}{\Gamma(\frac{M}{2}+1-n)}$
$\int d^{d}x_{1}\ldots d^{d}x_{n}~J(x_{1})\ldots J(x_{n})\int\mathcal{D}_{M}\phi~\phi(x_{1})\ldots\phi(x_{n})\left(1+\frac{2}{M}\frac{z}{\hbar}S[\phi]\right)^{\frac{M}{2}-n}\,.$ (72)
@@
$\left(1+\frac{2}{M}\frac{z}{\hbar}S[\phi]\right)^{\frac{M}{2}-n}=e^{\frac{z}{\hbar}S[\phi]}\exp\Bigg(\sum_{j=1}^{\infty}(-1)^{j}~\frac{2(2j-2)!!}{(j+1)!}\left(\frac{z}{\hbar}\frac{S[\phi]}{M}\right)^{j}~[jS[\phi]+(j+1)n]\Bigg)\,.$ (73)
@@
$\Omega[\mathcal{A}_{z},J]=\kappa_{M}~\left(\frac{\hbar M}{2z}\right)^{\frac{M}{2}}~\sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{z}{\hbar}\right)^{n}\left(\frac{2}{M}\right)^{n}\frac{\Gamma(\frac{M}{2}+1)}{\Gamma(\frac{M}{2}+1-n)}$
$\int d^{d}x_{1}\ldots d^{d}x_{n}~J(x_{1})\ldots J(x_{n})~\left\langle\phi(x_{1})\ldots\phi(x_{n})~e^{\sum_{j=1}^{\infty}(-1)^{j}~\frac{2(2j-2)!!}{(j+1)!}\left(\frac{z}{\hbar}\frac{S[\phi]}{M}\right)^{j}~[jS[\phi]+(j+1)n]}\right\rangle,$ (74)
where the expectation value, denoted $\langle\,\cdot\,\rangle$, is taken with respect to the weight $\exp\left(zS[\phi]/\hbar\right)$ with $S[\phi]$ being the renormalized action. Consider now the correlators:
$\left\langle\phi(x_{1})\ldots\phi(x_{n})~e^{\sum_{j=1}^{\infty}(-1)^{j}~\frac{2(2j-2)!!}{(j+1)!}\left(\frac{z}{\hbar}\frac{S[\phi]}{M}\right)^{j}~[jS[\phi]+(j+1)n]}\right\rangle=$
$=\left\langle\phi(x_{1})\ldots\phi(x_{n})\right\rangle+\sum_{j=1}^{\infty}\frac{c_{j}(M,n)}{M^{j}}\left\langle\phi_{1}(x_{1})\ldots\phi_{n}(x_{n})~S^{j}[\phi]\right\rangle,$ (75)
@@
$\lim_{M\to\infty}\left\langle\phi_{1}(x_{1})\ldots,\phi_{n}(x_{n})\frac{c_{j}(M,n)}{M^{j}}S^{j}[\phi]\right\rangle=0~~~\forall~j,$ (78)
@@
$\Omega[\mathcal{A}_{z},J]=$
$=\kappa_{M}~\left(\frac{\hbar M}{2z}\right)^{\frac{M}{2}}~\sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{z}{\hbar}\right)^{n}~\int d^{d}x_{1}\ldots d^{d}x_{n}~J(x_{1})\ldots J(x_{n})~\left\langle\phi(x_{1})\ldots\phi(x_{n})\right\rangle$
$=\kappa_{M}~\left(\frac{\hbar M}{2z}\right)^{\frac{M}{2}}\int\mathcal{D}\phi_{M}~e^{\frac{z}{\hbar}S[\phi]}\sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{z}{\hbar}\right)^{n}\int d^{d}x_{1}\ldots d^{d}x_{n}\,J(x_{1})\ldots J(x_{n})\phi(x_{1})\ldots\phi(x_{n})$
$=\kappa_{M}~\left(\frac{\hbar M}{2z}\right)^{\frac{M}{2}}\int\mathcal{D}\phi_{M}~e^{\frac{z}{\hbar}S[\phi]+\frac{z}{\hbar}\int d^{d}x~J(x)\phi(x)}\,.$ (79)
@@
It comes quite natural at this point a short remark on how the continuum limit can be possibly be considered in our numerical setup, in order to gain full consistency between simulation results and the present analytical derivation. Very simply we assume that in the symplectic quantization framework the continuum limit can be taken exactly as in ordinary lattice quantum field theory. One has to consider the limit of a vanishing lattice spacing, $a\rightarrow 0$, while tuning the bare parameters so as to hold a chosen renormalization group (RG) invariant observable fixed. In practice, the prescription to consider the continuum limit can be realized as follows: after having identified a given RG-invariant quantity $X$, one has to perform simulations at several values of the lattice spacings $a$, then adjusting the coupling(s) and the mass(es) so that the RG-invariant observable $X$ remains constant, thus ensuring a correct RG flow. The results of simulations can be then used to extrapolate smoothly the values of physical observables at $a=0$, i.e., in the continuum limit. Let us also notice that this procedure, while being necessary for any future use of the Symplectic Quantization approach to extract physical information for realistic theories, it is not particularly interesting for an asymptotically trivial theory such as $\lambda\phi^{4}$, where the RG-flow is to a Gaussian fixed point, which makes the continuum limit trivial.
@@
This is completely in agreement with the results of numerical simulations in the microcanonical ensemble, where the Hamiltonian dynamics of the free theory develops run-away solutions. Second, the result of our derivation in Eq. (80) shows us that this new “canonical Minkowskian measure” can be connected to the standard Feynman path integral by means of analytic continuation in the dimensionless parameter $z$. More investigations in this direction are actually in progress.
8 Conclusions and Perspectives